OpenStudy (anonymous):

if a gas occupies 10mL at 298K and 90kPa what volume does it occupy at STP

OpenStudy (wolfe8):

First use PV=nRT to find the number of moles. Then multiply with volume of gas at STP http://erhs.rcsnc.org/UserFiles/Servers/Server_4766394/Image/Donna%20Scofield/mole%20conversion%20map.bmp

OpenStudy (anonymous):

That will work. Can also use ratio of absolute temperatures and ratio of pressures V/Vo = (T/To)(Po/P)

OpenStudy (jfraser):

use the combined gas law:\[\frac{P_1V_1}{T_1} = \frac{P_2V_2}{T_2}\] and solve for the volume at STP